OpenStudy (anonymous):

Help please! How to find the derivative of x^2(3x-9)^3 by using the chain rule?

OpenStudy (tamtoan):

what is the formula for taking derivative of a product ?

OpenStudy (anonymous):

product rule is f'(g)+f(g')= \[dy/dx\]

OpenStudy (anonymous):

but I'm not using the product rule I'm using the chain which is \[h'(x)= f(g(x)) * g'(x)\]

OpenStudy (tamtoan):

f'(uv) ? ...if i remember correct ly...(uv)' = u'v + uv', isn't it right?

OpenStudy (anonymous):

Yurp that's the same product rule o.o

OpenStudy (tamtoan):

ok let your u = x^2 and v = (3x -9)^3 , can you try to get (uv)' using that formula ?

OpenStudy (anonymous):

I got \[2x(3x-9)^3+x^2(3(3x-9)^2\] haha hope that's right

OpenStudy (tamtoan):

v(x) = (3x - 9)^3 ..check your v'(x) again please.

OpenStudy (anonymous):

\[v'(x) = 3(3x-9)^2 *3?\] sorry I'm not good at math, that's what it says on my paper I just don't know how to get to it

OpenStudy (tamtoan):

i think is' 3(3x-9)^2 . 3 the second 3 is from taking derivative of 3x - 9 so you will have 9(3x-9)^2 ...now sub all in and expand them out to simplify

OpenStudy (tamtoan):

i forgot to suggest you factor first,,,probably don't need to expand them out, too much work

OpenStudy (anonymous):

OHHHHHHHHH I GET IT haha so \[x(3x-9)^2(2(3x-9) + 9x)\] for the factored ahhhh thanks you helped a ton XD that was the right answer that I wrote down

OpenStudy (tamtoan):

:) ok, hope you enjoy.